OpenStudy (anonymous):

When 0.422g of phosphorus is burned 0.967g of a white oxide is obtained. Determine the empirical formula of the oxide

OpenStudy (shikamaru11):

now what's that we gonna use the combustion analysis here too @Photon336 Sir???

OpenStudy (photon336):

@Shalante this is your type of problem.

OpenStudy (photon336):

i'm not sure myself @Shikamaru11 thoughts?

OpenStudy (rushwr):

Combustion is when u burn it in air soo we have to write the balanced equation for the reaction between phosphorus and O2

OpenStudy (shikamaru11):

Sir will be the equation is something like that?????????? \[P+ZnO \rightarrow ????\] hope i could think far

OpenStudy (rushwr):

Wait I dnt think we need to go for the equation. We can just assume that all of the phosphorus is now in the oxide.

OpenStudy (photon336):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Rushwr Combustion is when u burn it in air soo we have to write the balanced equation for the reaction between phosphorus and O2 \(\color{#0cbb34}{\text{End of Quote}}\) yeah, was thinking that too, BTW where did the zinc come from?

OpenStudy (rushwr):

Considering phosphorus as the limiting agent

OpenStudy (shikamaru11):

oh I got that in a second shot I think it mentioned white oxide but it's the obtant so my hypothesis failed :P

OpenStudy (shikamaru11):

hey my guess of PO4 lolxxxx :P

OpenStudy (rushwr):

we only need to find the empirical formula of the oxide right? So what if we consider that all of the phosphorus present are in the oxide now ? So the rest of the mass that is from 0.967 is oxygen SO the mass of oxygen in the oxide = 0.967 - 0.422 = 0.545g We can find the no. of moles of each separately Moles = mass/ molar mass moles of P = 0.422/ 31 = 0.014 moles Moles of Oxygen = 0.545/16 = 0.034 moles Now divide each by the smallest no. of moles  Now the ratio between P: O is 1:2.5 But we can't keep the ration in decimal numbers so to make it a whole number we multiply it by 2 so the new ratio would be 2:5 So the empirical formula would be P2O5